OpenStudy (anonymous):

Brad sold candy bars and cookies for a fundraiser at school. Candy bars sold for $3 and cookies sold for $5. He sold a total of 20 items and made $76. How many candy bars did Brad sell? A.12 B.14 C.16 D.18 need some help

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Candy bars=b cookies=c total of 20 items so b+c=20 $76 made so 3b+5c=76 solve for c in the first equation c=20-b plug that into second equation for c 3b+5(20-b)=76 now can you solve for b?

OpenStudy (anonymous):

ok thanks